OpenStudy (anonymous):

Need help Solving equation! 7(x+2) + 3 = -5(x-3) -1

OpenStudy (anonymous):

@Hero

pooja195 (pooja195):

Distribute 7(x+2)

OpenStudy (anonymous):

7x+14

pooja195 (pooja195):

7x+14+3=-5(x-3) -1 Distribute -5(x-3)

OpenStudy (anonymous):

-5x+15

pooja195 (pooja195):

7x+14+3=-5x+15-1 now combine like terms

OpenStudy (anonymous):

i got 12x=-9

OpenStudy (anonymous):

so -3/9?

pooja195 (pooja195):

Nope whats 14+3 ?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

oops i see my mistake

OpenStudy (anonymous):

give me a sec

pooja195 (pooja195):

hmm ok try again :)

OpenStudy (anonymous):

12x= -11

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Simplify both sides of the equation. 7x+14+3=−5x+15−1 Simplify 7x+14+3=−5x+15−1 7x+14+3=−5x+15+−1 (7x)+(14+3)=(−5x)+(15+−1)(Combine Like Terms) 7x+17=−5x+14 7x+17=−5x+14 Add 5x to both sides. 7x+17+5x=−5x+14+5x 12x+17=14 Subtract 17 from both sides. 12x+17−17=14−17 12x=−3 Divide both sides by 12.

OpenStudy (anonymous):

OK i got -4

OpenStudy (anonymous):

or would we leave it as a fraction